OpenStudy (anonymous):

What happen when u subtract 0

OpenStudy (anonymous):

\[0 - 0 = 0\] \[1 - 0 = 1\] \[2 - 0 = 2\] I think you understand that.

OpenStudy (compassionate):

You get @Qumecka 's smart score

OpenStudy (anonymous):

I got what I get, and what I get is what it is.

thomaster (thomaster):

Nothing happens when you subtract 0. Also this is posted in the wrong section :)